# Towards Agentic Intelligence for Materials Science

### v1

@@ 3.5.3 Multimodality
Recent advances in applying artificial intelligence to materials research have demonstrated the benefits of domain-specific language models, graph-based predictors, and image analysis frameworks 318, 250, 113. To effectively discover novel materials, agentic LLMs must process and understand multimodal data at each stage of the pipeline, not only as more inputs, but as a prerequisite for closing the loop, where each stage and component produces distinct modalities that must be jointly interpreted for decision-making 403, 371, 280. Most existing improvements are limited to single-model scenarios, where algorithms are designed for only a single type of data. However, scientific research often uses multiple types of input for representation, such as text reports, structural graphs, microscopy images, and spectroscopic signals. Each representation reveals the material’s properties from a different perspective. Integrating different types of data into a single analytical framework is challenging due to methodological and infrastructure issues. This section will explore these issues, first introducing the characteristics of multimodal scientific data, and then gradually delving into the various methods designed to integrate the data, with success defined in terms of cross-modal generalization, more reliable mechanistic inference, and improved closed-loop decision-making.
@@
Materials science research increasingly relies on multiple data sources, each revealing different aspects of material properties, thus requiring tailored data analysis methods. Using domain-specific language models, textual information from publications, patents, and experimental protocols has been put into order. MatSciBERT 145 showed that it is possible to pull out entities and relationships from scientific literature. The MatSci-NLP benchmark 329 set up standard tasks for synthesis action extraction, which made it possible to systematically compare and make further analysis on current methods.
@@
Image-based modalities provide further information at both mesoscopic and microscopic scales. There are big improvements in deep learning for microscopy, but a complete study 293 points out that some problems remains, such as failing to do 3D reconstruction, high reliance on synthetic data, and lacking enough annotated dataset. Another approach is through spectroscopic and diffraction data. Interpretable XRD analysis 397 shows that multi-phase quantification remain difficult, especially in the absence of simulation-based guidance. DiffractGPT 422 analyzes diffraction profiles in a novel way, treating them as sequence data, enabling autoregressive modeling methods to directly infer atomic structures.
@@
Recent workshops have enriched the research environment by providing multimodal datasets and benchmarks. The TDCM25 dataset 282 gathers crystalline materials that change with temperature, along with aligned structural, image, and textual characteristics, which helps with research on phase transitions. Similarly,  Takeda et al. 346 showed a multimodal fundamental dataset that integrates SELFIES molecular representations, DFT-calculated properties, and spectral characteristics to improve the accuracy of property predictions. NanoMINER 262 utilized language models to extract structured knowledge on nanomaterials, connecting unstructured literature with structured graphs. These improvements show that multimodal resources are becoming more comprehensive, but they also show that there is still no common metadata or ontologies that can be used with each other. Because of this, it is still hard to reproduce results across different modalities.
@@
Beyond crystalline and bulk materials, molecular discovery introduces additional representational challenges that require foundation models to be ‘multilingual’ across structural, physicochemical, and knowledge-based modalities. To accelerate scientific discovery, a foundation model must become “multilingual,” able to interpret the vast modalities of scientific data beyond text. This includes 2D 374 and 3D 215 structures described using geometric graphs, physicochemical properties (e.g., chirality, electrostatic potentials, and hydrophobicity), and dynamic data from molecular and quantum dynamics (MD/QD). It must also incorporate knowledge from literature 102 and knowledge bases 201, such as images, charts 428, and reaction pathways 374.
@@
Some recent work proposed joint molecule and language learning 103 and applied it to power conversion efficiency for organic photovoltaic (OPV) devices 255. However, most molecule language models tokenize at the level of atoms and/or they do not infuse synthesis considerations up front, which precludes matching tokens with desired functions and/or leads to AI-generated new molecular targets that are not readily synthesizable. To address these challenges, the recent highly innovative, function- and synthesis-aware modular Chemical Language Model (mCLM) 102 proposed a unified multimodal architecture and a suite of specialized encoders to tokenize each modality into a biologically and chemically meaningful language that can be reasoned over. mCLM incorporates function- and synthesis-related knowledge into the molecule discovery process a priori. mCLM generates new data on demand in the forward direction by an AI guided closed loop, where a human scientist simply inputs a few natural language phrases or sentences describing the desired functions of a molecule and receives the molecular structure with the best synthesizable functional modules in return.
@@
However, some additional challenges remain. The self-attention mechanism in the Transformer operates by allowing each token to aggregate information from other tokens based on semantic relevance 365. This formulation is rooted in natural language processing, where word embeddings capture semantic associations and attention primarily aggregates unary information across tokens. However, such a design may not align well with materials discovery in the physical world, where governing mechanisms are often determined by explicit pairwise or relational interactions rather than semantic similarity 128, 32. For example, in organic photovoltaic (OPV) materials, key performance-determining factors—such as donor–acceptor energy level alignment and balanced charge transport—are jointly determined by interacting molecular pairs 308, 53. This mismatch suggests that future research should explore attention mechanisms that explicitly encode binary or relational “mutual influence” between material entities, introducing inductive biases more consistent with physical interaction mechanisms 32, 128. Besides, while unified embedding-based multimodal models often improve predictive accuracy, they frequently provide limited mechanistic insight, raising concerns that learned representations may reflect correlations rather than causal or physically grounded relationships 313.
@@
Recent methodological advancements have concentrated on developing unified representation frameworks that integrate multiple modalities into a single representation space. MultiMat 248 presented a self-supervised foundation model that simultaneously encodes text, structural graphs, microscope pictures, and spectroscopic profiles, thereby augmenting transfer learning for subsequent property prediction. MatBind 245 built upon this by using contrastive learning aims to align density-of-states, crystal structures, and textual descriptions, achieving strong cross-modal retrieval. MatMCL 391 also enhanced the robustness of the model when dealing with data that only has some modalities, proving that multimodal information is redundant and complementary, and that the model can maintain prediction accuracy even when one or more modal information is missing.
@@
Other solutions change the way language models work, enabling them to handle non-textual data. For example, DiffractGPT 422 uses sequential modeling approaches to analyze diffraction spectra. This shows that methods originally designed for natural language can also be applied to crystallography. Hybrid systems that combine (GNNs, such as M3GNet 68) with LLMs further demonstrate how atomic representations can be linked to contextual knowledge in text. There are also new efforts to set benchmarks. MatVQA 113 establishes multimodal question answering using figures and text, whereas MaCBench 11 consolidates band structures, microscope images, and literature, facilitating a systematic assessment of general-purpose multimodal models in scientific endeavors. Fusion models like MATMMFuse 270 use multi-head attention to combine crystal graph embeddings with SciBERT textual encoders. This consistently makes it easier to predict formation energies and electrical characteristics. However, most current multimodal benchmarks emphasize recognition and cross-modal alignment, and only partially evaluate multi-step scientific reasoning or closed-loop decision-making—the regimes most relevant to autonomous discovery.
@@
These strategies substantially help bring together different signals, but there are still certain problems that need to be fixed. Unified embeddings can improve forecasts, but they don’t help us understand how things work very well. This raises the possibility that the captured relationship might show correlations instead of causal ties. Using high-dimensional modalities like microscope images increases computational and data-engineering costs, which makes them harder to access. Evaluation protocols are still fragmented. For example, MatSci-NLP 329 only covers text-based tasks, and there are no common benchmarks for multimodal reasoning or cross-modal generalization. TDCM25 282 and MaCBench 11 are two other efforts that show progress toward shared resources. However, there is currently no consensus in the sector about evaluation suites. Future research directions encompass the integration of physical invariances into fusion architectures and the establishment of standardized multimodal datasets featuring interoperable ontologies, both of which are essential for ensuring that multimodal AI significantly advances scientific discovery and can be evaluated by its contribution to end-to-end discovery outcomes (e.g., experiment selection, validation throughput, and constraint satisfaction), not only predictive metrics.
@@ 3.6 Explainability
Explainable AI is becoming a key component of AI-based applications in materials science. This change is happening because the field increasingly requires models that go beyond forecasting to provide mechanistic insight, error diagnosis, and trustworthy decision-making 426. This is considerably more important if we use autonomous agentic LLMs in real-world tests, where explanations can serve as verification artifacts that gate high-stakes experimental actions. There are no universally accepted criteria for evaluating interpretations among groups. Alvarez-Melis and Jaakkola 16 came up with three dimensions: explicitness, faithfulness, and stability. Explicitness examines the clarity and comprehensibility of the explanations, faithfulness assesses whether the relevance score accurately reflects the true determinants of the model output, and stability assesses the consistency of explanations derived from similar inputs. Other research also makes a distinction between plausibility and faithfulness. Plausibility refers to consistency with human thinking, while faithfulness refers to adhering to the internal thought process of the model 10. In a comparable framework, Lipton 220 delineated three tiers of model interpretability: simulatability (the entirety of the model is comprehensible to humans), decomposability (the individual components are interpretable), and algorithmic transparency (the learning process is comprehensible and reproducible).
@@ 3.6.1 Sparse and Closed-form Models
In materials science, few explainable models are more prominent than those that generate sparse closed-form representations between input features and target properties. The SISSO family of algorithms268, 401, 251, symbolic regression methods applied for perovskite catalysis 388 scientific machine learning with sparse regression for stability equations 252, and dimension-constrained symbolic regression for band gap prediction 370. Unlike post-hoc explainers, these methods directly return human-readable algebraic formulas that may be considered candidate scientific laws or descriptors. This makes them particularly appealing in domains like materials science, where physical interpretability and mechanism discovery are paramount. However, these methods typically operate on curated feature sets and may struggle with raw, high-dimensional modalities, motivating complementary neural explainers discussed next.
@@ 3.6.2 Attention and Graph Explainers
Whereas sparse regression methods provide intrinsic interpretability by expressing structure–property relations in closed-form descriptors268, most state-of-the-art AI models in materials science are deep neural networks, in particular GNNs and transformer-based architectures426, 375. These models achieve strong predictive performance across a wide range of tasks but are often criticized as “black boxes”, since their internal decision logic is difficult to inspect directly. In response, recent work has introduced both embedded and post-hoc explanation strategies to make such models more transparent by highlighting which atoms, bonds, compositions, or spectral features drive individual predictions426, 375.
@@
Representative examples include CrabNet 369, a compositionally constrained attention network whose attention mechanisms surface influential elemental constituents of a compound; GANN 319, a graph attention networks for 2D materials property prediction that localizes importance to specific atoms and bonds; CrysXPP 91, which trains crystal GNNs under property-sensitive explanatory paths to connect predictions with interpretable substructures; and spectroscopy-focused interpreters that generate attribution maps of spectra to identify peaks that contribute most to the output 193. Together, these approaches move beyond raw accuracy toward more transparent materials models that can be interrogated by domain experts, even though important challenges remain in ensuring that the resulting explanations are faithful, stable, and aligned with underlying physical mechanisms.
@@
Attention-based and graph-based explainers have strengths in their capacity to provide local, instance-particular reasons, like determining which atoms, bonds or spectral peaks guided the most prediction375, 9. This increases the physical validity of their measurements, as highlighted elements tend to conform to established chemical intuition. Nevertheless, faithfulness is not always guaranteed: attention weights may not correspond to the model’s actual decision logic, raising the question of whether highlighted features are causally responsible or only correlated161.
@@
Stability still remains a hurdle, because inputs perturbed in small amounts or random seeds can modify the highlighted substructures or spectra9. There are multiple approaches, including integrated gradients337. It appears that the durability of these explanations in materials models has been studied, and the proposed robustness is explored by perturbation-based tests9. Thus, overall, attention- and graph-based explanation approaches are highly desirable for supplying domain-plausible, instance-level insights. They are great for alignment with human chemical reasoning (plausibility/validity) but suffer from challenges in faithfulness and stability. These limitations motivate hybrid approaches that have a combination approach: attention using perturbation tests or physics-informed constraints, leading into the next subsection on physics-informed interpretability.
@@ 3.6.3 Physics-informed Interpretability
A third group of solutions provides intrinsic interpretability by embedding known physical principles directly into the model design. Physics-informed methods constrain predictions to physically consistent families, which builds trust and usefulness in materials research426. They don’t rely on explanations that come after the fact. Symmetry-equivariant networks, constraint-based architectures, or the strict implementation of conservation and dimensional norms can all help with this34. For example, equivariant graph neural networks that take into account the translational, rotational, and permutational symmetries of crystalline systems 34; models that are limited by physical invariants like energy conservation or stress–strain relationships 238.
The fundamental benefit of physics-informed interpretability is that it focuses on physical validity: predictions follow the known structure of the issue space without making up correlations 411. Physics-informed interpretability enhances faithfulness, stability, and safety by confining models to physically valid actions, acting as critical feasibility guards in closed-loop discovery. But these limits can make it hard to be flexible and generalize when the laws of physics are complicated or not fully understood. This means that we need hybrid approaches that combine clear physical rules with data-driven learning to make strong, agentic pipelines.

### v2

@@ 3.5.3 Multimodality
Recent advances in applying artificial intelligence to materials research have demonstrated the benefits of domain-specific language models, graph-based predictors, and image analysis frameworks 320, 252, 114. To effectively discover novel materials, agentic LLMs must process and understand multimodal data at each stage of the pipeline, not only as more inputs, but as a prerequisite for closing the loop, where each stage and component produces distinct modalities that must be jointly interpreted for decision-making 406, 374, 282. Most existing improvements are limited to single-model scenarios, where algorithms are designed for only a single type of data. However, scientific research often uses multiple types of input for representation, such as text reports, structural graphs, microscopy images, and spectroscopic signals. Each representation reveals the material’s properties from a different perspective. Integrating different types of data into a single analytical framework is challenging due to methodological and infrastructure issues. This section will explore these issues, first introducing the characteristics of multimodal scientific data, and then gradually delving into the various methods designed to integrate the data, with success defined in terms of cross-modal generalization, more reliable mechanistic inference, and improved closed-loop decision-making.
@@
Materials science research increasingly relies on multiple data sources, each revealing different aspects of material properties, thus requiring tailored data analysis methods. Using domain-specific language models, textual information from publications, patents, and experimental protocols has been put into order. MatSciBERT 146 showed that it is possible to pull out entities and relationships from scientific literature. The MatSci-NLP benchmark 331 set up standard tasks for synthesis action extraction, which made it possible to systematically compare and make further analysis on current methods.
@@
Image-based modalities provide further information at both mesoscopic and microscopic scales. There are big improvements in deep learning for microscopy, but a complete study 295 points out that some problems remains, such as failing to do 3D reconstruction, high reliance on synthetic data, and lacking enough annotated dataset. Another approach is through spectroscopic and diffraction data. Interpretable XRD analysis 400 shows that multi-phase quantification remain difficult, especially in the absence of simulation-based guidance. DiffractGPT 425 analyzes diffraction profiles in a novel way, treating them as sequence data, enabling autoregressive modeling methods to directly infer atomic structures.
@@
Recent workshops have enriched the research environment by providing multimodal datasets and benchmarks. The TDCM25 dataset 284 gathers crystalline materials that change with temperature, along with aligned structural, image, and textual characteristics, which helps with research on phase transitions. Similarly,  Takeda et al. 348 showed a multimodal fundamental dataset that integrates SELFIES molecular representations, DFT-calculated properties, and spectral characteristics to improve the accuracy of property predictions. NanoMINER 264 utilized language models to extract structured knowledge on nanomaterials, connecting unstructured literature with structured graphs. These improvements show that multimodal resources are becoming more comprehensive, but they also show that there is still no common metadata or ontologies that can be used with each other. Because of this, it is still hard to reproduce results across different modalities.
@@
Beyond crystalline and bulk materials, molecular discovery introduces additional representational challenges that require foundation models to be ‘multilingual’ across structural, physicochemical, and knowledge-based modalities. To accelerate scientific discovery, a foundation model must become “multilingual,” able to interpret the vast modalities of scientific data beyond text. This includes 2D 377 and 3D 216 structures described using geometric graphs, physicochemical properties (e.g., chirality, electrostatic potentials, and hydrophobicity), and dynamic data from molecular and quantum dynamics (MD/QD). It must also incorporate knowledge from literature 103 and knowledge bases 202, such as images, charts 431, and reaction pathways 377.
@@
Some recent work proposed joint molecule and language learning 104 and applied it to power conversion efficiency for organic photovoltaic (OPV) devices 257. However, most molecule language models tokenize at the level of atoms and/or they do not infuse synthesis considerations up front, which precludes matching tokens with desired functions and/or leads to AI-generated new molecular targets that are not readily synthesizable. To address these challenges, the recent highly innovative, function- and synthesis-aware modular Chemical Language Model (mCLM) 103 proposed a unified multimodal architecture and a suite of specialized encoders to tokenize each modality into a biologically and chemically meaningful language that can be reasoned over. mCLM incorporates function- and synthesis-related knowledge into the molecule discovery process a priori. mCLM generates new data on demand in the forward direction by an AI guided closed loop, where a human scientist simply inputs a few natural language phrases or sentences describing the desired functions of a molecule and receives the molecular structure with the best synthesizable functional modules in return.
@@
However, some additional challenges remain. The self-attention mechanism in the Transformer operates by allowing each token to aggregate information from other tokens based on semantic relevance 368. This formulation is rooted in natural language processing, where word embeddings capture semantic associations and attention primarily aggregates unary information across tokens. However, such a design may not align well with materials discovery in the physical world, where governing mechanisms are often determined by explicit pairwise or relational interactions rather than semantic similarity 129, 33. For example, in organic photovoltaic (OPV) materials, key performance-determining factors—such as donor–acceptor energy level alignment and balanced charge transport—are jointly determined by interacting molecular pairs 310, 54. This mismatch suggests that future research should explore attention mechanisms that explicitly encode binary or relational “mutual influence” between material entities, introducing inductive biases more consistent with physical interaction mechanisms 33, 129. Besides, while unified embedding-based multimodal models often improve predictive accuracy, they frequently provide limited mechanistic insight, raising concerns that learned representations may reflect correlations rather than causal or physically grounded relationships 315.
@@
Recent methodological advancements have concentrated on developing unified representation frameworks that integrate multiple modalities into a single representation space. MultiMat 250 presented a self-supervised foundation model that simultaneously encodes text, structural graphs, microscope pictures, and spectroscopic profiles, thereby augmenting transfer learning for subsequent property prediction. MatBind 247 built upon this by using contrastive learning aims to align density-of-states, crystal structures, and textual descriptions, achieving strong cross-modal retrieval. MatMCL 394 also enhanced the robustness of the model when dealing with data that only has some modalities, proving that multimodal information is redundant and complementary, and that the model can maintain prediction accuracy even when one or more modal information is missing.
@@
Other solutions change the way language models work, enabling them to handle non-textual data. For example, DiffractGPT 425 uses sequential modeling approaches to analyze diffraction spectra. This shows that methods originally designed for natural language can also be applied to crystallography. Hybrid systems that combine (GNNs, such as M3GNet 69) with LLMs further demonstrate how atomic representations can be linked to contextual knowledge in text. There are also new efforts to set benchmarks. MatVQA 114 establishes multimodal question answering using figures and text, whereas MaCBench 11 consolidates band structures, microscope images, and literature, facilitating a systematic assessment of general-purpose multimodal models in scientific endeavors. Fusion models like MATMMFuse 272 use multi-head attention to combine crystal graph embeddings with SciBERT textual encoders. This consistently makes it easier to predict formation energies and electrical characteristics. However, most current multimodal benchmarks emphasize recognition and cross-modal alignment and only partially evaluate multi-step scientific reasoning or closed-loop decision-making—the regimes most relevant to autonomous discovery.
@@
These strategies substantially help bring together different signals, but there are still certain problems that need to be fixed. Unified embeddings can improve forecasts, but they don’t help us understand how things work very well. This raises the possibility that the captured relationship might show correlations instead of causal ties. Using high-dimensional modalities like microscope images increases computational and data-engineering costs, which makes them harder to access. Evaluation protocols are still fragmented. For example, MatSci-NLP 331 only covers text-based tasks, and there are no common benchmarks for multimodal reasoning or cross-modal generalization. TDCM25 284 and MaCBench 11 are two other efforts that show progress toward shared resources. However, there is currently no consensus in the sector about evaluation suites. Future research directions encompass the integration of physical invariances into fusion architectures and the establishment of standardized multimodal datasets featuring interoperable ontologies, both of which are essential for ensuring that multimodal AI significantly advances scientific discovery and can be evaluated by its contribution to end-to-end discovery outcomes (e.g., experiment selection, validation throughput, and constraint satisfaction), not only predictive metrics.
@@ 3.6 Explainability
Explainable AI is becoming a key component of AI-based applications in materials science. This change is happening because the field increasingly requires models that go beyond forecasting to provide mechanistic insight, error diagnosis, and trustworthy decision-making 429. This is considerably more important if we use autonomous agentic LLMs in real-world tests, where explanations can serve as verification artifacts that gate high-stakes experimental actions. There are no universally accepted criteria for evaluating interpretations among groups. Alvarez-Melis and Jaakkola 16 came up with three dimensions: explicitness, faithfulness, and stability. Explicitness examines the clarity and comprehensibility of the explanations, faithfulness assesses whether the relevance score accurately reflects the true determinants of the model output, and stability assesses the consistency of explanations derived from similar inputs. Other research also makes a distinction between plausibility and faithfulness. Plausibility refers to consistency with human thinking, while faithfulness refers to adhering to the internal thought process of the model 10. In a comparable framework, Lipton 222 delineated three tiers of model interpretability: simulatability (the entirety of the model is comprehensible to humans), decomposability (the individual components are interpretable), and algorithmic transparency (the learning process is comprehensible and reproducible).
@@ 3.6.1 Sparse and Closed-form Models
In materials science, few explainable models are more prominent than those that generate sparse closed-form representations between input features and target properties. The SISSO family of algorithms270, 404, 253, symbolic regression methods applied for perovskite catalysis 391 scientific machine learning with sparse regression for stability equations 254, and dimension-cons-trained symbolic regression for band gap prediction 373. Unlike post-hoc explainers, these methods directly return human-readable algebraic formulas that may be considered candidate scientific laws or descriptors. This makes them particularly appealing in domains like materials science, where physical interpretability and mechanism discovery are paramount. However, these methods typically operate on curated feature sets and may struggle with raw, high-dimensional modalities, motivating complementary neural explainers discussed next.
@@ 3.6.2 Attention and Graph Explainers
Whereas sparse regression methods provide intrinsic interpretability by expressing structure-property relations in closed-form descriptors270, most state-of-the-art AI models in materials science are deep neural networks, in particular GNNs and transformer-based architectures429, 378. These models achieve strong predictive performance across a wide range of tasks but are often criticized as “black boxes”, since their internal decision logic is difficult to inspect directly. In response, recent work has introduced both embedded and post-hoc explanation strategies to make such models more transparent by highlighting which atoms, bonds, compositions, or spectral features drive individual predictions429, 378.
@@
Representative examples include CrabNet 372, a compositionally constrained attention network whose attention mechanisms surface influential elemental constituents of a compound; GANN 321, a graph attention networks for 2D materials property prediction that localizes importance to specific atoms and bonds; CrysXPP 92, which trains crystal GNNs under property-sensitive explanatory paths to connect predictions with interpretable substructures; and spectro-scopy-focused interpreters that generate attribution maps of spectra to identify peaks that contribute most to the output 194. Together, these approaches move beyond raw accuracy toward more transparent materials models that can be interrogated by domain experts, even though important challenges remain in ensuring that the resulting explanations are faithful, stable, and aligned with underlying physical mechanisms.
@@
Attention-based and graph-based explainers have strengths in their capacity to provide local, instance-particular reasons, like determining which atoms, bonds or spectral peaks guided the most prediction378, 9. This increases the physical validity of their measurements, as highlighted elements tend to conform to established chemical intuition. Nevertheless, faithfulness is not always guaranteed: attention weights may not correspond to the model’s actual decision logic, raising the question of whether highlighted features are causally responsible or only correlated162.
@@
Stability still remains a hurdle, because inputs perturbed in small amounts or random seeds can modify the highlighted substructures or spectra9. There are multiple approaches, including integrated gradients339. It appears that the durability of these explanations in materials models has been studied, and the proposed robustness is explored by perturbation-based tests9. Thus, overall, attention- and graph-based explanation approaches are highly desirable for supplying domain-plausible, instance-level insights. They are great for alignment with human chemical reasoning (plausibility/validity) but suffer from challenges in faithfulness and stability. These limitations motivate hybrid approaches that have a combination approach: attention using perturbation tests or physics-informed constraints, leading into the next subsection on physics-informed interpretability.
@@ 3.6.3 Physics-informed Interpretability
A third group of solutions provides intrinsic interpretability by embedding known physical principles directly into the model design. Physics-informed methods constrain predictions to physically consistent families, which builds trust and usefulness in materials research429. They don’t rely on explanations that come after the fact. Symmetry-equivariant networks, constraint-based architectures, or the strict implementation of conservation and dimensional norms can all help with this35. For example, equivariant graph neural networks that take into account the translational, rotational, and permutational symmetries of crystalline systems 35; models that are limited by physical invariants like energy conservation or stress–strain relationships 240.
The fundamental benefit of physics-informed interpretability is that it focuses on physical validity: predictions follow the known structure of the issue space without making up correlations 414. Physics-informed interpretability enhances faithfulness, stability, and safety by confining models to physically valid actions, acting as critical feasibility guards in closed-loop discovery. But these limits can make it hard to be flexible and generalize when the laws of physics are complicated or not fully understood. This means that we need hybrid approaches that combine clear physical rules with data-driven learning to make strong, agentic pipelines.